Solve the linear partial differential equation (D⁴+D'⁴)=0

1
Expert's answer
2021-11-29T15:42:04-0500

Solution;

Given;

"D^4+D'^4=0"

The auxiliary equation is;

"m^4+1=0"

"m^4=-1=e^{i\u03c0}" (By Euler's formula)

By De Moivres theorem;

"m^4=e^{i\u03c0+2n\u03c0}"

"m=e^{iπ(\frac{2n+1}{4})}" In which n=0,1,2,3....

When;

n=0;

"m=e^{i\fracπ4}=cos\fracπ4+isin\fracπ4=\frac{1}{\sqrt2}+i\frac{1}{\sqrt2}"

n=1;

"m=e^{i(\frac{3π}{4})}=cos\frac{3π}{4}+isin\frac{3π}{4}=\frac{-1}{\sqrt2}+i\frac{1}{\sqrt2}"

n=2;

"m=e^{i(\frac{5π}{4})}=cos\frac{5π}{4}+i sin\frac{5π}{4}=\frac{-1}{\sqrt2}-i\frac{1}{\sqrt2}"

n=3;

"m=e^{i (\frac{7π}{4})}=cos\frac{7π}{4}+i sin\frac{7π}{4}=\frac{1}{\sqrt2}-i\frac{1}{\sqrt2}"

Therefore,you can see that the roots are;

"m=\frac{1}{\sqrt2}_-^+i\frac{1}{\sqrt2}" and "-\frac{1}{\sqrt2}_-^+\frac{1}{\sqrt2}"

Therefore,the complementary solution is;

"C.F=e^{\frac{1}{\sqrt2}x}[c_1cos\frac{1}{\sqrt2}x+c_2sin\frac{1}{\sqrt2}x]+e^{\frac{-1}{\sqrt2}}[c_3cos\frac{1}{\sqrt2}x+c_4sin\frac{1}{\sqrt2}x]"